OpenStudy (anonymous):

SAT PRACTICE QUESTION: The sum of three numbers is 6. Each number is increased by 20 and the new numbers are multiplied by 10. What is the sum of the resulting numbers? Could you help me set up an equation?

OpenStudy (kainui):

You have 3 numbers, let's call them a, b, and c. What do we know? a+b+c=6. Can you try making more equations like this based off the information given? Take your best guess we'll help you out.

OpenStudy (anonymous):

\(a+b+c=6\) \[10(a+20+b+20+c+20)=?\]

OpenStudy (anonymous):

10a + 200 + 10b + 200 + 10 c + 200 = 10a + 10b + 10c + 600

OpenStudy (anonymous):

you got it?

OpenStudy (anonymous):

It says the answer is 660. Why is that?

OpenStudy (anonymous):

make 10a+10b+10c=x-600 solve for a+b+c

OpenStudy (anonymous):

10a + 10b + 10c = x - 600 a + b + c = x/10 - 6

OpenStudy (anonymous):

Ok yes, I got it: 10(a + 20) + 10(b + 20) + 10(c + 20) = required sum 10a + 200 + 10b + 200 + 10c + 200 = 600 + 10(a + b + c) Since a + b + c = 6; the sum = 600 + 60 = 660